- Nakano, Enomoto, Yokozawa, Yokogawa, Tomaru, Ushiba Guiguo : Some work for Y arm alignment.
Transition from morning work (*:responsible person)
- If you open chamber or flange, you must put a label on the door/flange for "when/who" oped/closed door, and leave "current status".
- COMM: SR alignment will be proceeded in the middle or late of next week
Report of work (*:responsible person)
- IMC: ASC loop was closed. Lock lasted all the night.
- COMM: Beam is hitting on the center of PR2.
- PR2: We saw the big pitch difference on Oplev as before.
- ITMY : IP matrix tuning was done.
- Needs more time for damping. ITMY will be ready next week for SR alignment.
- MIR: Yokozawa, Kimura, Tomaru checked the surface of ETMY mirror. It was found that the surface was very dirty by eye and Tcam.
- It is not sure for the cleaning.
- WAB ETMX: Completed and moved in the safe mode. Needs a small remaining work for ITMY.
Work plan (*:responsible person)
- PR2 and PR3: we will open doors.
- Type B: needs tuning for BS frequency. GEO phone is something wrong.
- DGS: X end rack maintenance will finish this Friday or beginning of next week.